# Rigidity Criterion for Certain Calabi–Yau families

### v1

@@ Abstract.
We prove a new rigidity criterion for families of polarized Calabi-Yau manifolds. Motivated by known non-rigid examples, we conjecture that a family over a quasi-projective curve is rigid if it admits a smooth compactification whose singular fiber has only isolated singularities. We verify this conjecture for singularities with a concentrated mixed Hodge spectrum– class including ordinary double points and cusps. The proof combines an analysis of the vanishing cycle exact sequence and limiting mixed Hodge structure with a tensor-product decomposition of the associated variation of Hodge structures.
@@ 1. Introduction
Let $f\colon(\mathcal{X},\mathcal{L})\to S$ be a flat family of $n$-dimensional polarized smooth Calabi–Yau manifolds over a quasiprojective curve $S$. Let $\overline{S}$ be a projective smooth compactification of $S$ with a boundary point $0\in\overline{S}\setminus S$. Suppose that the family extends to a flat family $\overline{\mathcal{X}}\to\overline{S}$ such that the total space $\overline{\mathcal{X}}$ is smooth and the singular fiber $X_{0}=f^{-1}(0)$ has only isolated singularities.
@@
Then by Deligne’s theorem, each irreducible factor of $\mathbb{V}$ decompose as tensor product of VHS pulled back from $S$ and $B$ respectively. Are these tensor-product decompositions motivic? More precisely, can one find some product family $(g,h)\colon\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to S^{\prime}\times B^{\prime}$ ($S^{\prime}$ and $B^{\prime}$ are finite coverings of $S$ and $B$) such that after base changing these tensor-products can be realized as irreducible factors of $R^{\cdot}g_{*}\mathbb{C}_{\mathcal{Y}_{1}}\otimes R^{\cdot}h_{*}\mathbb{C}_{\mathcal{Y}_{2}}$?
@@ 3.3. Period integral approach to ODP
When the isolated singularities of $X_{0}$ are ordinary double points, there is a differential geometric approach by period integrals. The argument here is from the introduction part of [RT09], see [Wan97, Corollary 3.5] for a proof via Clemens–Schmid exact sequence. Here the nontriviality of the vanishing cycles is from the estimates in geometric analysis of [CGPY23, Lemma 4.3].
@@ 4. Tensor-product decomposition of limiting MHS
By Proposition 3.5 or [Wan97, Corollary 3.5], the monodromy operator $T$ on $\mathrm{H}_{\mathrm{lim}}^{3}(X_{t})$ satisfies that $N\neq 0$ and $T^{ss}=Id$. So the boundary point corresponding to the singular fiber $X_{0}$ is at infinite distance with respect to the Hodge metric on the period domain of $\mathbb{V}_{1}$. Since the two period domains corresponding to $\mathbb{V}_{1}$ and $\mathbb{V}_{2}$ are both Hermitian symmetric domainis, the Hodge metric is the same as the Weil–Petersson metric up to a constant multiple. Therefore, the Weil–Petersson metric on $S$ is complete along the smoothing for the singular fiber $X_{0}$. This contradicts to C.-L. Wang’s result [Wan97, Proposition 2.3].

### v2

@@ Abstract.
We prove a new rigidity criterion for families of polarized Calabi-Yau manifolds. Motivated by known non-rigid examples, we conjecture that a family over a quasi-projective curve is rigid if it admits a smooth compactification for which one of the singular fiber has only isolated singularities. We verify this conjecture for singularities with a concentrated mixed Hodge spectrum—a class including ordinary double points and cusps. The proof combines an analysis of the vanishing cycle exact sequence and limiting mixed Hodge structure with a tensor-product decomposition of the associated variation of Hodge structures.
@@ 1. Introduction
Let $f\colon(\mathcal{X},\mathcal{L})\to S$ be a flat family of $n$-dimensional polarized smooth Calabi–Yau manifolds over a quasiprojective curve $S$. Let $\overline{S}$ be a projective smooth compactification of $S$ with a boundary point $0\in\overline{S}\setminus S$. Suppose that the family extends to a flat family $\overline{\mathcal{X}}\to\overline{S}$ such that the total space $\overline{\mathcal{X}}$ is smooth near $f^{-1}(0)$ and the singular fiber $X_{0}=f^{-1}(0)$ has only isolated singularities.
@@
Then by Deligne’s theorem, each irreducible factor of $\mathbb{V}$ decompose as tensor product of VHS pulled back from $S$ and $B$ respectively. Are these tensor-product decompositions motivic? More precisely, can one find some product family $(g,h)\colon\mathcal{Y}_{1}\times\mathcal{Y}_{2}\to S^{\prime}\times B^{\prime}$ ($S^{\prime}$ and $B^{\prime}$ are finite coverings of $S$ and $B$) and a correspondence between the base change of $f\colon\mathcal{X}\to S\times B$ and the product family $(g,h)$ such that these tensor-products can be realized as irreducible factors of $R^{\cdot}g_{*}\mathbb{C}_{\mathcal{Y}_{1}}\otimes R^{\cdot}h_{*}\mathbb{C}_{\mathcal{Y}_{2}}$?
@@ 3.3. Period integral approach to ODP
When the isolated singularities of $X_{0}$ are ordinary double points. There is a differential geometric approach by period integrals. The argument here is from the introduction part of [RT09], see [Wan97, Corollary 3.5] for a proof via Clemens–Schmid exact sequence. Here the nontriviality of the vanishing cycles is from the estimates in geometric analysis of [CGPY23, Lemma 4.3].
@@ 4. Tensor-product decomposition of limiting MHS
By Proposition 3.5 or [Wan03, Theorem 1.5(1)], the monodromy operator $T$ on $\mathrm{H}_{\mathrm{lim}}^{3}(X_{t})$ satisfies that $N\neq 0$ and $T^{ss}=Id$. So the boundary point corresponding to the singular fiber $X_{0}$ is at infinite distance with respect to the Hodge metric on the period domain of $\mathbb{V}_{1}$. Since the two period domains corresponding to $\mathbb{V}_{1}$ and $\mathbb{V}_{2}$ are both Hermitian symmetric domainis, the Hodge metric is the same as the Weil–Petersson metric up to a constant multiple. Therefore, the Weil–Petersson metric on $S$ is complete along the smoothing for the singular fiber $X_{0}$. This contradicts to C.-L. Wang’s result [Wan97, Proposition 2.3].